OpenStudy (anonymous):

Recycled cds, inc, offers a choice of 5 used cds for $20, with each additional cd costing $6. Write a cost function 5 or more cds. What will the ocst of buying 14 cds be? Let x represent the number of cds over 5 C(x)=?

OpenStudy (anonymous):

@ajspeller thank you soooo muuch!!!!!

OpenStudy (anonymous):

@ajspeller what happened?

OpenStudy (ajspeller):

c(x) = 6(x-5) + 20 c(14) = 6(14-5) + 20 c(14) = 6(9) + 20 c(14) = 54 + 20 c(14) = 74 c(5) = 6(5-5) + 20 c(5) = 6(0) + 20 c(5) = 0 + 20 c(5) = 20

OpenStudy (anonymous):

@ajspeller ok thank you so much!